# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ A.3 Proofs Section 6
(Wasserstein Distance under Approximated Stochastic Linear Operations) Let $p_{{w}}\in{\mathcal{P}}({\mathbb{R}}^{(n+1)\times m})$ and $q_{{w}}\in{\mathcal{P}}({\mathbb{R}}^{(n+1)\times m})$ be weight distributions, and denote $p_{L_{{w}}({{z}})}=L_{{{w}}}({{z}})\#p_{{w}}$ and $q_{L_{{w}}({{z}})}=L_{{{w}}}({{z}})\#q_{{w}}$. Then, for probability distributions $p,q\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$ it holds that
${\mathbb{W}}_{\rho}({\mathbb{E}}_{{{z}}\sim p}[p_{L_{{w}}({{z}})}],{\mathbb{E}}_{\tilde{{{z}}}\sim q}[q_{L_{{w}}(\tilde{{{z}}})}])\leq{\mathbb{E}}_{{{W}}\sim p_{{w}}}[\|{{W}}\|^{\rho}]^{\frac{1}{\rho}}{\mathbb{W}}_{\rho}(p,q)+{\mathbb{E}}_{\tilde{{{z}}}\sim q}[\|\tilde{{{z}}}\|^{\rho}]^{\frac{1}{\rho}}{\mathbb{W}}_{\rho}(p_{{W}},q_{{{W}}})+{\mathbb{W}}_{\rho}(p_{{b}},q_{{{b}}}).$
Proof  By the triangular inequality:
${\mathbb{W}}_{\rho}({\mathbb{E}}_{{{z}}\sim p}[p_{L_{{w}}({{z}})}],{\mathbb{E}}_{\tilde{{{z}}}\sim q}[q_{L_{{{w}}}(\tilde{{{z}}})}])$ $\leq\underbrace{{\mathbb{W}}_{\rho}({\mathbb{E}}_{{{z}}\sim p}[p_{L_{{w}}({{z}})}],{\mathbb{E}}_{\tilde{{{z}}}\sim q}[p_{L_{{{w}}}(\tilde{{{z}}})}])}_{\text{(a)}}+\underbrace{{\mathbb{W}}_{\rho}({\mathbb{E}}_{{{z}}\sim q}[p_{L_{{w}}({{z}})}],{\mathbb{E}}_{\tilde{{{z}}}\sim q}[q_{L_{{{w}}}(\tilde{{{z}}})}])}_{\text{(b)}}.$
For term (a), we derived upper bound ${\mathbb{E}}_{{{W}}\sim p_{{w}}}[\|{{W}}\|^{\rho}]^{\frac{1}{\rho}}{\mathbb{W}}_{\rho}(p,q)$ in Proposition 25. For term (b), we can apply the Minkowski inequality to obtain:
${\mathbb{W}}_{\rho}({\mathbb{E}}_{{{z}}\sim q}[p_{L_{{w}}({{x}})}],{\mathbb{E}}_{\tilde{{{z}}}\sim q}[q_{L_{{{w}}}(\tilde{{{z}}})}])$
$\qquad\qquad=\inf_{\gamma\in\Gamma(p_{{w}},q_{{w}},q)}{\mathbb{E}}_{(({{W}},{{b}}),(\tilde{{{W}}},\tilde{{{b}}}),\tilde{{{z}}})\sim\gamma}[\|{{W}}\tilde{{{z}}}-\tilde{{{W}}}\tilde{{{z}}}+{{b}}-\tilde{{{b}}}\|^{\rho}]^{\frac{1}{\rho}}$
$\qquad\qquad\leq\inf_{\gamma\in\Gamma(p_{{w}},q_{{w}},q)}{\mathbb{E}}_{(({{W}},{{b}}),(\tilde{{{W}}},\tilde{{{b}}}),\tilde{{{z}}})\sim\gamma}[\|({{W}}-\tilde{{{W}}})\tilde{{{z}}}\|^{\rho}]^{\frac{1}{\rho}}+{\mathbb{E}}_{(({{W}},{{b}}),(\tilde{{{W}}},\tilde{{{b}}}),\tilde{{{z}}})\sim\gamma}[{{b}}-\tilde{{{b}}}\|^{\rho}]^{\frac{1}{\rho}}$
$\qquad\qquad=\underbrace{\inf_{\gamma_{{b}}\in\Gamma(p_{{b}},q_{{b}})}{\mathbb{E}}_{({{b}},\tilde{{{b}}})\sim\gamma_{{b}}}[{{b}}-\tilde{{{b}}}\|^{\rho}]^{\frac{1}{\rho}}}_{\text{(b.1)}}+\underbrace{\inf_{\gamma\in\Gamma(p_{{W}},q_{{W}},q)}{\mathbb{E}}_{(({{W}},{{b}}),(\tilde{{{W}}},\tilde{{{b}}}),\tilde{{{z}}})\sim\gamma}[\|({{W}}-\tilde{{{W}}})\tilde{{{z}}}\|^{\rho}]^{\frac{1}{\rho}}.}_{\text{(b.2)}}$
Here, $p_{{b}},p_{{W}}$ and $q_{{b}},q_{{W}}$ denote the marginals of $p_{{w}}$ and $q_{{w}}$ w.r.t. ${{b}}$ and ${{W}}$, respectively. Note that term (b.1) is equivalent to ${\mathbb{W}}_{\rho}(p_{{b}},q_{{b}})$. Using the triangle inequality, we can upper bound term (b.2) as
$\inf_{\gamma\in\Gamma(p_{{W}},q_{{W}},q)}{\mathbb{E}}_{(({{W}},{{b}}),(\tilde{{{W}}},\tilde{{{b}}}),\tilde{{{z}}})\sim\gamma}[\|({{W}}-\tilde{{{W}}})\tilde{{{z}}}\|^{\rho}]$ $\leq\inf_{\gamma\in\Gamma(p_{{W}},q_{{W}},q)}{\mathbb{E}}_{({{W}},\tilde{{{W}}},\tilde{{{b}}}),\tilde{{{z}}})\sim\gamma}[\|{{W}}-\tilde{{{W}}}\|^{\rho}\|\tilde{{{z}}}\|^{\rho}]$
$\leq{\mathbb{E}}_{\tilde{{{z}}}\sim q}[\|\tilde{{{z}}}\|^{\rho}]\inf_{\gamma\in\Gamma(p_{{W}},q_{{W}})}{\mathbb{E}}_{({{W}},\tilde{{{W}}},\tilde{{{z}}})\sim\gamma}[\|{{W}}-\tilde{{{W}}}\|^{\rho}]$
$\leq{\mathbb{E}}_{\tilde{{{z}}}\sim q}[\|\tilde{{{z}}}\|^{\rho}]{\mathbb{W}}_{\rho}^{\rho}(p_{{W}},q_{{W}}).$
Combining the above bounds, we obtain the inequality as in Eqn. (26).
Corollary 27 (of Proposition 8 on Lipschitz Transformations)
For $p\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$ and signature locations ${\mathcal{C}}=\{{{c}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{n}$, let ${\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}$ be the local Lipschitz constant of piecewise continuous function ${\sigma}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ in region ${\mathcal{R}}_{i}$, then for $\rho\in\mathbb{N}$
${\mathbb{W}}_{\rho}^{\rho}({\sigma}\#p,{\sigma}\#{\Delta}_{\mathcal{C}}\#p)\leq\sum_{i=1}^{N}{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}^{\rho}{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-{{c}}_{i}\|^{\rho}\mid{{z}}\in{\mathcal{R}}_{i}].$
Proof  According to Theorem 1 in (Ambrogioni et al., 2018) we have that
Then, as by definition it hold that $\|{\sigma}({{z}})-{\sigma}({{c}}_{i})\|\leq{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}\|{{z}}-{{c}}_{i}\|$, $\forall{{z}},{{c}}_{i}\in{\mathcal{R}}_{i}$, we have:
Note also that, for ${\mathcal{L}}_{\sigma}$ the global Lipschitz constant, i.e., ${\mathcal{L}}_{\sigma}=\max_{i\in\{1,\ldots,N\}}{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}$, it follows that
$\sum_{i=1}^{N}{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}^{\rho}{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-{{c}}_{i}\|^{\rho}\mid{{z}}\in{\mathcal{R}}_{i}]\leq{\mathcal{L}}_{{\sigma}}^{\rho}\sum_{i=1}^{N}{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-{{c}}_{i}\|^{\rho}\mid{{z}}\in{\mathcal{R}}_{i}]={\mathcal{L}}_{\sigma}^{\rho}{\mathbb{W}}^{\rho}_{\rho}(p,q).$ (32)
Proposition 28 (Triangle Inequality for Wasserstein Distance with Mappings)
For $\rho\in\{1,2\}$ and probability distributions $p,q\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$ and mapping $h:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$, it holds that
${\mathbb{W}}_{\rho}(p,h\#q)\leq{\mathbb{W}}_{\rho}(p,q)+{\mathbb{W}}_{\rho}(q,h\#q)$ (33)
Proof  The proof follows directly from the fact that ${\mathbb{W}}_{\rho}$ satisfies the axioms of a distance, including the triangle inequality.
Theorem 13: Wasserstein Distance between the Distribution of a SNN and its GMM Approximation Over a Point
Proof  We prove Theorem 13 via induction and generalize the result to the $\rho$-Wasserstein distance with $\rho\in\{1,2\}$. To do so, we first derive the base case of the induction. For $k\in\{1,\ldots,K\}$, we have that:
${\mathbb{W}}_{\rho}(p_{nn({{x}}),k+1},q_{nn({{x}}),k+1})$
$\qquad={\mathbb{W}}_{\rho}({\mathbb{E}}_{{{{z}}}_{k}\sim p_{nn({{x}}),k}}[p_{L_{{{w}}_{k}}({\sigma}({{{z}}}_{k}))}],{\mathbb{E}}_{\tilde{{{{z}}}}_{k}\sim{R}_{M}(q_{nn({{x}}),k})}[p_{L_{{{w}}_{k}}({\sigma}({\Delta}_{{\mathcal{C}}_{k}}(\tilde{{{{z}}}}_{k})))}])$ Eqns. (4) & (13)
$\qquad\leq{\mathbb{E}}_{{{W}}_{k}\sim p_{{{w}}_{k}}}[\|{{W}}\|^{\rho}]^{\frac{1}{\rho}}{\mathbb{W}}_{\rho}({\sigma}\#p_{nn({{x}}),k},{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({{x}}),k}))$ Prop. 25
$\qquad\leq\hat{\mathcal{S}}_{\rho,{{w}},k}{\mathbb{W}}_{\rho}({\sigma}\#p_{nn({{x}}),k},{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({{x}}),k}))$ Eqn. (15)
$\qquad\leq\hat{\mathcal{S}}_{\rho,{{w}},k}\bigl{[}\underbrace{{\mathbb{W}}_{\rho}({\sigma}\#p_{nn({{x}}),k},{\sigma}\#{R}_{M}(q_{nn({{x}}),k}))}_{\text{(I)}}$
$\qquad\qquad+\underbrace{{\mathbb{W}}_{\rho}({\sigma}\#{R}_{M}(q_{nn({{x}}),k}),{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({{x}}),k}))}_{\text{(II)}}\bigr{]}$ Prop. 28
where, for (I),
${\mathbb{W}}_{\rho}({\sigma}\#p_{nn({{x}}),k},{\sigma}\#{R}_{M}(q_{nn({{x}}),k}))$
$\qquad\leq{\mathcal{L}}_{\sigma}{\mathbb{W}}_{\rho}(p_{nn({{x}}),k},{R}_{M}(q_{nn({{x}}),k}))$ Corol. 27
$\qquad\leq{\mathcal{L}}_{\sigma}\left[{\mathbb{W}}_{\rho}(p_{nn({{x}}),k},q_{nn({{x}}),k})+{\mathbb{W}}_{\rho}(q_{nn({{x}}),k},{R}_{M}(q_{nn({{x}}),k}))\right]$ Prop. 28
$\qquad\leq{\mathcal{L}}_{\sigma}\left[{\mathbb{W}}_{\rho}(p_{nn({{x}}),k},q_{nn({{x}}),k})+\hat{{\mathbb{W}}}_{\rho,{R},k}\right],$ Eqn. (17)
with $L_{\sigma}$ is the global Lipschitz constant of $\sigma$, and for (II),
${\mathbb{W}}_{\rho}({\sigma}\#{R}_{M}(q_{nn({{x}}),k}),{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({{x}}),k}))\leq\hat{{\mathbb{W}}}_{\rho,{\Delta},k}$ Eqn. (16)
Hence, if ${\mathbb{W}}_{\rho}(p_{nn({{x}}),k},q_{nn({{x}}),k})\leq\hat{{\mathbb{W}}}_{\rho,k}$, then
${\mathbb{W}}_{\rho}(p_{nn({{x}}),k+1},q_{nn({{x}}),k+1})$ $\leq\hat{\mathcal{S}}_{\rho,{{w}},k}\left[{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,k}+{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,{R},k}+\hat{{\mathbb{W}}}_{\rho,{\Delta},k}\right]\coloneqq\hat{{\mathbb{W}}}_{\rho,k+1}$
${\mathbb{W}}_{\rho}(p_{nn({{x}})},q_{nn({{x}})})\leq\hat{{\mathbb{W}}}_{\rho,K+1}.$
Proposition 16: Wasserstein Distance Between the Distribution of a SNN and its GMM Approximation Over Set of Points
Proof  We prove this Proposition via induction and generalize the result to the $\rho$-Wasserstein distance with $\rho\in\{1,2\}$. We first derive the base case. For $\rho\in\{1,2\}$ and $k\in\{1,\ldots,K\}$,
${\mathbb{W}}_{\rho}(p_{nn({\mathcal{X}}),k+1},q_{nn({\mathcal{X}}),k+1})$
$\qquad={\mathbb{W}}_{\rho}({\mathbb{E}}_{{{{z}}}_{k}\sim p_{nn({\mathcal{X}}),k}}[p_{L^{D}_{{{w}}_{k}}({\sigma}({{{z}}}_{k}))}],{\mathbb{E}}_{\tilde{{{{z}}}}_{k}\sim{R}_{M}(q_{nn({\mathcal{X}}),k})}[p_{L^{D}_{{{w}}_{k}}({\sigma}({\Delta}_{{\mathcal{C}}_{k}}(\tilde{{{{z}}}}_{k})))}])$ Eqns. (21) & (22)
[…]
where, following the same steps as in the proof of Theorem 13,
${\mathbb{W}}_{\rho}({\sigma}\#p_{nn({\mathcal{X}}),k},{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{R}_{M}(q_{nn({\mathcal{X}}),k}))\leq{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,k}+{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,{R},k}+\hat{{\mathbb{W}}}_{\rho,{\Delta},k}.$
Further, since by the properties of the spectral norm $\|{{\bar{1}}}_{D}\otimes{{W}}\|^{\rho}=\|{{\bar{1}}}_{D}\|^{\rho}\|{{W}}\|^{\rho}$ and $\|{{\bar{1}}}_{D}\|^{\rho}=D^{\frac{\rho}{2}}$, we have
${\mathbb{E}}_{{{W}}_{k}\sim p_{{{w}}_{k}}}[\|{{\bar{1}}}_{D}\otimes{{W}}\|^{\rho}]^{\frac{1}{\rho}}$ $=(D^{\frac{\rho}{2}}{\mathbb{E}}_{{{W}}_{k}\sim p_{{{w}}_{k}}}[\|{{W}}\|^{\rho}])^{\frac{1}{\rho}},$
$\leq\hat{\mathcal{S}}_{\rho,{{w}},k}$ Eqn. (23)
Hence, if ${\mathbb{W}}_{\rho}(p_{nn({\mathcal{X}}),k},q_{nn({\mathcal{X}}),k})\leq\hat{{\mathbb{W}}}_{\rho,k}$, then

### v2

@@ Appendix E Compression for Dropout-Induced Bernoulli Mixtures
Let $d_{out}\in{\mathcal{D}}_{N\cdot 2^{n}}({\mathbb{R}}^{n})$ and $\bar{d}_{out}\in{\mathcal{D}}_{M}({\mathbb{R}}^{n})$ be as defined in Eqn (44) and Eqn (45), respectively. Then,
${\mathbb{W}}_{2}^{2}(d_{out},\bar{d}_{out})\leq\sum_{i\in{\mathcal{I}}^{c}}(1-\theta)\|{{c}}_{i}\|^{2}$ (46)
Proof  Note that we can write
$d_{out}=\sum_{i=1}^{N}{{{\pi}}^{(i)}}p_{i},$
where $p_{i}=\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\odot{{c}}_{i}}$, and
$\bar{d}_{out}=\sum_{i\in{\mathcal{I}}}{{{\pi}}^{(i)}}p_{i}+\sum_{i\in{\mathcal{I}}^{c}}{{{\pi}}^{(i)}}\delta_{{{c}}_{i}}.$
Hence, according to Lemma 31,
${\mathbb{W}}_{2}^{2}(d_{out},\bar{d}_{out})=\sum_{i\in{\mathcal{I}}}{{{\pi}}^{(i)}}{\mathbb{W}}_{2}^{2}\left(p_{i},p_{i}\right)+\sum_{i\in{\mathcal{I}}^{c}}{{{\pi}}^{(i)}}{\mathbb{W}}_{2}^{2}\left(p_{i},\delta_{{{c}}_{i}}\right)=\sum_{i\in{\mathcal{I}}^{c}}{{{\pi}}^{(i)}}{\mathbb{W}}_{2}^{2}\left(p_{i},\delta_{{{c}}_{i}}\right)$
To compute ${\mathbb{W}}_{2}^{2}\left(p_{i},\delta_{{{c}}_{i}}\right)$, we observe that the optimal transport plan moves all probability mass from $p_{i}$ to ${{c}}_{i}$. Thus,
${\mathbb{W}}_{2}^{2}\left(p_{i},\delta_{{{c}}_{i}}\right)$ $={\mathbb{E}}_{{{b}}\sim d_{b}}[\|{{b}}\odot{{c}}_{i}-{{c}}_{i}\|^{2}]={\mathbb{E}}_{{{b}}\sim d_{b}}\left[\sum_{j=1}^{n}\left({{{b}}^{(j)}}{{{c}}_{i}^{{(j)}}}-{{{c}}_{i}^{{(j)}}}\right)^{2}\right]=\sum_{j=1}^{n}{{{c}}_{i}^{{(j)}}}^{2}{\mathbb{E}}_{\beta\sim p_{\beta}}\left[(\beta-1)^{2}\right],$
where $p_{\beta}=\theta\delta_{1}+(1-\theta)\delta_{0}$. Thus
${\mathbb{E}}_{\beta\sim p_{\beta}}\left[(\beta-1)^{2}\right]=1-{\mathbb{E}}_{\beta\sim p_{\beta}}\left[\beta^{2}-2\beta\right]=1-{\mathbb{E}}_{\beta\sim p_{\beta}}\left[\beta\right]=1-\theta.$